The bicycle wheel makes 5 revolutions. Determine how far the bicycle travels in feet. The diameter of the wheel is 45in. Use 3.14 for pie
Answer by stanbon(75887)

The bicycle wheel makes 5 revolutions. Determine how far the bicycle travels in feet. The diameter of the wheel is 45in. Use 3.14 for pie
----
Circumference of the wheel = (pi)d = 45*3.14 = 141.3 inches
----
Distance traveled in inches: 5*141.3 = 706.5 inches
---
Distance traveled in feet: 706.5/12 = 58.875 feet
====================================================
Cheers,
Stan H.
=========